OpenStudy (anonymous):

Calculate the number of moles of gas in 1500. km3 of air at a temperature of 15 ∘C and a pressure of 792.0torr .

OpenStudy (abb0t):

Use the ideal gas law formula to solve for \(\sf \color{red}{n}\). Now, the ideal gas law formula is: \(\sf \color{orange}{PV = nRT}\). As you might remember from algebra, to rearrange the formula. HINT: that means divide.

OpenStudy (abb0t):

P = pressure [atm] V = volume [L] n = mole R = gas constant; which is either, 0.0821 \(\frac{liter~\times~atm}{mol \times K}\) OR: 8.3145 \(\frac{J}{mol \times K}\) OR: 8.2057 \(\frac{m3 \times atm}{mol\times K}\) and finally, T = temperature [K] As you will see, based on the gas constant, \(\sf \color{blue}{R}\) that you use, you will need to convert some units to match those of the constant.